OpenStudy (anonymous):

find the exact value of the expression for the given value of 0. Do not use a calculator. Sin(30) if 0 = Pi/6

OpenStudy (anonymous):

I'm guessing that you were trying to write θ instead of 0. So first we plug in the value of θ into sin(3θ). What you are looking for is the sin(3*π/6) or sin(π/2). Using the unit circle (which I am assuming you know how to use) we can see that sin(π/2) is 1.

OpenStudy (anonymous):

Stevex, thanks for the reply, yes I meant theta. I am not sure how to use the unit circle well. I just googled it earlier and have been trying to understand it. How did you come to sin(3*π/6)?

OpenStudy (noelgreco):

I'm having a bit of trouble with the question. sin(30) = 0.5 pi/6 radians is equivalent to 30 degrees. sin(pi/6) = 0.5

OpenStudy (anonymous):

Steve is right, but I don't know how to come to that conclusion.

OpenStudy (anonymous):

i have to add the question make no sense what does "if \(\theta=\frac{\pi}{6}\) " have to do with \(\sin(30)\) ?

OpenStudy (anonymous):

This is one of the problems straight from the program.

OpenStudy (anonymous):

Sorry I just got back. This is a pretty good explanation of the unit circle: http://www.mathsisfun.com/geometry/unit-circle.html Also you can just convert the radians into degrees if you want (remember that π = 180 degrees). And then the sin(90 degrees) would just be 1. Hope I helped!

OpenStudy (anonymous):

I'm still kind of confused on how you got the original answer. " first we plug in the value of θ into sin(3θ). What you are looking for is the sin(3*π/6)" if sin 30 = .5 as per my calculator, and as per http://library.thinkquest.org/20991/alg2/trig.html (which is just a different version of the unit circle).

OpenStudy (anonymous):

if it is \(\sin(3\theta)\) and \(\theta=\frac{\pi}{6}\) then \[\sin(3\times \frac{\pi}{6})=\sin(\frac{\pi}{2})=1\]

OpenStudy (anonymous):

got it now, thanks everyone!